Two small sphere spaced 20.0 cm apart have equal charge. How many excess electrons must be present on each sphere if the magnitude of the force of repulsion between them is 4.57x10-21 N?



1
Expert's answer
2022-05-17T17:21:52-0400

"F=\frac{kn^2e^2}{r^2},\implies"

"n=\sqrt\frac Fk\cdot\frac re=891."